# inkscapelatex=false

### v2

@@ Abstract
Sequential activation of place-tuned neurons in an animal during navigation is typically interpreted as reflecting the sequence of input from adjacent positions along the trajectory. More recent theories about such place cells suggest sequences arise from abstract cognitive objectives like planning. Here, we propose a mechanistic and parsimonious interpretation to complement these ideas: hippocampal sequences arise from intrinsic recurrent circuitry that propagates activity without readily available input, acting as a temporal memory buffer for extremely sparse inputs. We implement a minimal sequence generator inspired by neurobiology and pair it with an actor–critic learner for egocentric visual navigation. Our agent reliably solves a continuous maze without explicit geometric cues, with performance depending on the length of the recurrent sequence. Crucially, the model outperforms LSTM cores under sparse input conditions (16 channels, $\sim 2.5\%$ activity), but not under dense input, revealing a strong interaction between representational sparsity and memory architecture. In contrast to LSTM agents, hidden sequence units develop localized place fields, distance-dependent spatial kernels, and task-dependent remapping, while inputs orthogonalize and spatial information increases across layers. These phenomena align with neurobiological data and are causal to performance. Together, our results show that sparse input synergizes with sequence-generating dynamics, providing both a mechanistic account of place cell sequences in the mammalian hippocampus and a simple inductive bias for reinforcement learning based on sparse egocentric inputs in navigation tasks.
Keywords Hippocampus $\cdot$ Visual navigation $\cdot$ Place cell $\cdot$ Theta rhythm $\cdot$ Reinforcement learning
@@ 1 Introduction
The hippocampus is a brain region that is essential for episodic memory and spatial navigation. Damage impairs memory formation in humans (Scoville and Milner, 2000), and rodent recordings revealed place cells tuned to specific locations (O’Keefe and Dostrovsky, 1971). These findings support the idea that spatial anchors structure episodic memory (Aronowitz and Nadel, 2023), reinforced by observations of replay sequences correlated to planning (Foster and Wilson, 2007; Kay et al., 2020).
@@
Recent computational models have recast these findings through modern machine learning. Successor representations interpret place cells as predictive states (Stachenfeld et al., 2017; Mattar and Daw, 2018); reservoir models emphasize pre-existing dynamics (Leibold, 2020); probabilistic approaches infer place fields as latent states from sequential sensory experiences(Raju et al., 2024); and self-supervised methods learn spatial tuning from movement statistics (Wang et al., 2024). While these approaches reproduce place-like activity and even sequential patterns, they rarely address explicitly where hippocampal sequences originate.
@@
We propose a parsimonious account: hippocampal theta sequences, that are experimentally observed during navigation Foster and Wilson (2007), arise from intrinsic recurrent circuitry Fenton (2024) in CA3 that can propagate activity over long timescales in the absence of input (Fig. 1.) The CA3 sequence generator receives sparse inputs from dentate gyrus (DG) to yield localized spatial codes in CA3 that support navigation (Leibold, 2020, 2022). The sparse-input regime is not incidental but central: it reflects ecological contraints of navigation often guided by only a few reliable landmarks amid abundant noisy sensory input; the biology of DG granule cells, which fire at extremely low mean rates; and the computational advantages of high-capacity, low-interference codes that promote high memory capacity, compositionality and generalization.
@@
From a machine learning perspective, our architecture resonates ideas from dynamical systems models for long-range memory such as LSTMs Hochreiter and Schmidhuber (1997) and echo state machines Jaeger and Haas (2004). More recently, state-space models (SSMs) and linear RNNs have been suggested that preserve information by expanding inputs into high-dimensional representations before compressing them with shallow nonlinear readouts (Fu et al., 2022; Gu et al., 2020, 2021). Our model extends these principles with a DG sparsification stage that generates a low-activity code optimized for the CA3 recurrent network, whose intrinsic sequence dynamics are particularly effective under sparse input conditions, offering a rich set of features for downstream policy learning from compact visual features.
@@
To test the hypothesis that sparse inputs to a long-horizon sequence-generating dynamical system are sufficient to give rise to a representation allowing successful navigation, we implement an agent with a DG-like sparsification module, a recurrent sequence generator (CA3 proxy), and an actor–critic learner for egocentric navigation. We show that sequence generation and sparse input synergize, outperforming LSTMs of comparable size in the sparse-input regime, while LSTMs remain competitive under dense input. Moreover, place-field like neural tuning, DG orthogonalization, and task-dependent remapping emerge naturally during training.
@@
Figure 1: Theta sequences. A Illustration of theta sequences observed in rodent hippocampus. In each theta cycle, R=3 neurons are activated and the activation propagates in a sequence of $\ell$=5 neurons over L=3 theta cycles (cf. eq. 2). B Current baseline understanding of the theta sequences, driven by sequential inputs despite recurrent connections in CA3. Left: spatial tunings of the sequentially activated cells. C Intrinsic theta sequence hypothesis, a parsimonious account where the recurrent connections support generating long horizon sequential activity without sequential external inputs.
@@ 2 Methods
Figure 2: Model summary. A Virtual environment ($19\times 19$ tiles) were constructed with walls randomly placed on 15 % of the tiles. Wall layouts for kept fixed for repeated trials, with an invisible goal (red star) near to the bottom right unless otherwise mentioned. At each episode, the agent was initially placed at a random location at least 5 tiles away from the goal. B The agent receives a first person perspective vision input that is processed via a visual encoder (ResNet, matching the SOTA in deepmind lab environment (Espeholt et al., 2018), pretrained and fixed in our experiments; cf. A3). C These output was linearly transformed to F features, and then sparsified using batch normalization and high thresholding ($\tau=2.43$), such that the percentage of activation ($\sim 2.5\%$) matches the sparse activity of neurons in the hippocampal DG that project to CA3 area. The modules in white are hard coded or fixed during training. The modules in blue (fully connected layers, FC) are trained.
@@ 2.1 Virtual Environment
We consider an agent navigating a continuous environment with sparse obstacles and uniform visual textures. The 3-d virtual environment was simulated using Deepmind lab (Beattie et al., 2016) with 60 fps, each action is repeated for 8 frames until the next observation. Visual observations are captured at half the wall height, with FOV of $90\degree$ and resolution of $92\times 76$ pixels in RGB channels. The map has $19\times 19$ squares (excluding the confining walls) with obstacle walls randomly placed on 15 % of the tiles. The wall layouts were kept fixed for repeated episodes unless otherwise stated. The environment was designed such that spatial relations between locations cannot be trivially inferred from the similarity of their corresponding visual features and that it allows multiple possible trajectories from a given location to the target.
In each episode, the agent was initially placed in a random location at least 5 units from the goal. An episode is finished when the agent reaches the goal location or a maximum of 7200 frames (with action repeat=8 corresponding to a maximum of 900 action steps). For the map displayed in Figure 2, the reward was placed near the bottom right corner and then moved to the lower left corner to test generalization. Since our navigation task does not require complex actions, we reduced the action space used in IMPALA (Espeholt et al., 2018), to facilitate training speed (Table A2).
@@ 2.2 Visual processing
The visual input is intended to mimic visual cortical preprocessing, extracting general-purpose visual features for the hippocampus. We pretrained a ResNet encoder (He et al., 2015) in combination with an LSTM core as in Espeholt et al. (2018) and kept it fixed when training our hippocampus model. We also tested a variant using the layer 2 output from ResNet pretrained on Imagenet. The model produces similar performance (Fig. A1).
@@ 2.3 Hippocampus Model
The main cortical input to the hippocampus is routed through the dentate gyrus (DG) (Amaral et al., 2007). DG activity is characteristically sparse, with only a small fraction of granule cells active in a given environment (Henze et al., 2000; Leutgeb and Leutgeb, 2007). We model the DG as a linear projection of visual features followed by batch normalization (running estimates of mean and variance, retention rate $0.95$ per minibatch) and a high activation threshold to match the sparsity of activity ($\sim 2.5\%$).
@@
We model CA3 as a linear RNN shift register that propagates inputs as theta sequences (Fig. 1A) and, for this paper, we keep it fixed to isolate the effects of intrinsic sequence generation from effects potentially induced by recurrent plasticity. The DG provides input $u_{t}\in\mathbb{R}^{F}$ over $F$ features. Each feature is assigned a dedicated prewired sequence of length $\ell$, so the CA3 state is $X_{t}\in\mathbb{R}^{F\ell}$. Motivated by hippocampal theta sequences (Dragoi and Buzsáki, 2006; Foster and Wilson, 2007; Leibold, 2020), we parameterize the total number of sequence units as $\ell=L+(R-1)$, where $L$ is the number of theta cycles and $R$ the number of active units per cycle.
@@
and the full CA3 dynamics
@@ 2.4 Actor Critic Network
The activity of hippocampal units is fed into a decoder module with two linear/fully connected (FC) layers with ReLU activation. The actions and value are computed as the linear readout from the decoder module. We trained the agent with a standard advantage actor–critic objective (policy-gradient + value-baseline + entropy regularization) as implemented in Sample Factory (Petrenko et al., 2020; Espeholt et al., 2018).
Figure 3: Training performance A Agents with different sequence length L and repetition number R. Line and shaded area are mean and s.e.m. across 6 random seeds. Len: number of frames to reach the goal location. Stable success: the rate of having 100 consecutive successful episodes. Metrics were Gaussian smoothed, $\sigma=6\times 10^{6}$ frames. B The best performing agent with L=64 and R=8 across all seeds was tested for its transfer learning. New Rew.: new reward location at the lower left corner. new map: a randomly generated map with the same statistics. block path: new walls are added to block paths to the reward, while the previous blocking walls are removed. C and D Performance of agents with different recurrent modules and inputs. C: sparse input. D: dense input where batch-normalization and high thresholding was removed. CA3: our CA3 model with L=64 and R=8. RandRNN: randomly initialized fixed RNN of the same state size. SSM_LegS: fixed SSM HiPPO-LegS from Gu et al. (2020) with the same state size. LSTM: trainable LSTM with matching number of parameters.
@@ 3.1 Behavioral Performance
Training the naive agent (L=64, R=8) on one fixed reward location in the maze exhibits robust performance measures after about 350 million frames (Fig. 3A). Thus our model effectively maintains information about sparse inputs in its trajectories of the RNN-like dynamics across time like in a reservoir (Jaeger and Haas, 2004). This view is further supported by our simulations with reduced sequence length showing inferior performance (Fig. 3 A). In the most extreme case ($L=1,R=1$) without sequences, essentially a pure feedforward architecture corresponding to a brain with DG output bypassing CA3, the agent did not achieve robust behavior. While agents with reduced sequence lengths ($L=1,16,R=8$) still can express reasonable success rates and trajectory lengths, their behavior is considerably more unstable as evidenced by the low fraction of consecutive successes (Fig. 3A). Transfer learning for new reward location and blocked paths, however, requires only about 50 million frames, and transfer learning to a new map requires about 150 million frames. These indicate that the agent was able to form a generalizing representation of the map and task (Fig. 3B).
@@
Crucially, this advantage is confined to the sparse-input regime. Under dense input, LSTMs perform better—consistent with prior reports (Hessel et al., 2019)—and our CA3 module performs worse than HiPPO-LegS and random RNNs (Fig. 3D). We also tested other variants from Gu et al. (2020) and they performed similar to HiPPO-LegS (Tab. 1, Fig. A3. These results highlight a specific synergy between sparse representations and intrinsic sequence generator.
Figure 4: Evolution of occupancy over the course of learning. Color represents a normalized measure of time the agent spent at a location. Mean head directions at a location were visualized in arrows. From epoch 5 onwards, the agent has a preference to reach the bottom right corner (independent of the random starting location) and proceed to the goal location (red star) from there.
@@ 3.2 Occupancy
We divided the training into 6 epochs and evaluated the behavior and learned representations at these checkpoints. The agent develops a stable trajectory after 4 to 5 epochs. Between Epoch 4 and Epoch 5, the agent learned to get around the obstacle in the upper part of the right edge and started to spend more time at the lower right corner before reaching the reward site (Fig. 4). The agent appears to develop a strategy of visiting locations with salient input/landmarks and converging from different starting locations to the habituated paths, even after switching to a new reward location. This goes in line with the typical strategies employed by animals in familiar enviroments Gibson et al. (2013). In comparison, the LSTM agent under dense sensory input has fewer converging trajectories before reaching the reward (Fig. A8, likely due to the goal information being more readily available in the visual input, thereby implementing a strategy more related to visual search.
@@
Figure 5: Left: spatial tuning of DG and CA3 units from epoch 6. Pixel coordinates correspond to environment (black crosses correspond to walls). Each row shows the CA3 units ordered by their positions in the activity sequence. We selected 4 out of the 16 feature sequences for visualization. Spatial Information (SI): bits per time step. Right: spatial tuning of randomly selected Decoder layer 1 units from epoch 6.
Figure 6: Correlation (color code) of the DG activity maps evoked by the 16 (rows and columns) visual features. Throughout training, the DG input to CA3 became increasingly orthogonal.

### v3

@@ Abstract
Sequential firing of hippocampal place cells is often attributed to sequential sensory drive along a trajectory, and has also been attributed to planning and other cognitive functions. Here, we propose a mechanistic and parsimonious interpretation to complement these ideas: hippocampal sequences arise from intrinsic recurrent circuitry that propagates transient input over long horizons, acting as a temporal memory buffer that is especially useful when reliable sensory evidence is sparse. We implement this idea with a minimal sequence generator inspired by neurobiology and pair it with an actor–critic learner for egocentric visual navigation. Our agent reliably solves a continuous maze without explicit geometric cues, with performance depending on the length of the recurrent sequence. Crucially, the model outperforms LSTM cores under sparse input conditions (16 channels, $\sim 2.5\%$ activity), but not under dense input, revealing a strong interaction between representational sparsity and memory architecture. Through learning, units develop localized place fields, distance-dependent spatial kernels, and task-dependent remapping, while inputs to the sequence generator orthogonalize and spatial information increases across layers. These phenomena align with neurobiological data and are causal to performance. Together, our results show that sparse input synergizes with sequence-generating dynamics, providing both a mechanistic account of place cell sequences in the mammalian hippocampus and a simple inductive bias for reinforcement learning based on sparse egocentric inputs in navigation tasks. Code: https://github.com/xiaoxionglin/SF_hipposlam
@@ 1 Introduction
Hippocampal place cells track the animal’s location during navigation (O’Keefe & Dostrovsky, 1971) and they fire in sequence reflecting the behavioral order of the place fields (Foster & Wilson, 2007). Spatial locations are thereby thought to serve as anchors for episodic memories (Aronowitz & Nadel, 2023), a view reinforced by observations of “look-ahead” sequence replay linked to trajectory planning (Foster & Wilson, 2007; Kay et al., 2020). On the other hand, hippocampal neurons are also found to fire at successive moments not explainable by location alone (Eichenbaum, 2014), suggesting the place cell sequences could reflect timing rather than spatial input.
@@
The intertwined spatial and temporal representations in hippocampus has been touched upon in many recent computational models. Successor representations interpret place cells as predictive states (Stachenfeld et al., 2017; Mattar & Daw, 2018); reservoir models emphasize pre-existing dynamics in shaping place cell sequences (Leibold, 2020); probabilistic approaches model place cells as latent states inferred from successive inputs (Raju et al., 2024); and self-supervised methods refine spatial tuning by exploiting temporal smoothness of the trajectory (Wang et al., 2024). While these approaches reproduce place-like activity and even sequential patterns, they rarely address explicitly where hippocampal sequences originate. Many bottom-up studies instead emphasize that structured representations emerge from circuit dynamics and physiological constraints, suggesting mechanistic models are needed to link anatomy, dynamics, and representational patterns (Lin et al., 2023; Buzsáki & Tingley, 2018; Schaeffer et al., 2022).
@@
We propose a parsimonious account directly addressing this question: hippocampal sequences arise from intrinsic recurrent circuitry in CA3 that can propagate activity over long timescales in the absence of input (Fig. 1A-C). The CA3 sequence generator receives sparse inputs from dentate gyrus (DG) to yield localized spatial codes in CA3 that support navigation (Leibold, 2020; 2022). The sparse-input regime is not incidental but central: it reflects the ecological reality that navigation is often guided by only a few reliable landmarks amid abundant sensory noise; the biology of DG granule cells, which fire at extremely low rates; and the computational advantages of high-capacity, low-interference codes that promote compositionality and generalization.
@@
This mechanism mirrors recent key ideas from machine learning. State-space models and structured linear RNNs preserve long-range information by expanding inputs into a high-dimensional temporal feature space before compressing them via shallow nonlinear readouts (Fu et al., 2022; Gu et al., 2020; 2021). Our model resonates with this principle: DG sparsification provides a low-activity code that is sustained and expanded by intrinsic recurrence, offering a rich set of features for downstream policy learning.
@@
To test this hypothesis, we implement an agent with a DG-like sparsification module, a recurrent sequence generator (CA3 proxy), and an actor–critic learner for egocentric navigation. We show that sequence generation and sparse input synergize, outperforming LSTMs of comparable size in the sparse-input regime, while LSTMs remain competitive under dense input. Moreover, place fields, DG orthogonalization, and task-dependent remapping emerge naturally during training.
@@ 2 Methods
Figure 1: Model summary. A: Illustration of theta sequences observed in rodent hippocampus. In each theta cycle, $R=3$ neurons are activated and the activation propagates over $L=4$ theta cycles in a sequence of $\ell=L+R-1=6$ neurons. B: The theta sequences are thought to be driven by sequential inputs despite the recurrent connections in hippocampus. C: The recurrent connections could support generating long horizon sequential activity without sequential external inputs. D: A virtual environment ($19\times 19$ tiles) was constructed using DeepMind Lab with walls randomly placed on 15 % of the tiles. Wall layouts are kept fixed across episodes, with an invisible reward near the bottom right. E: The agent receives a first person perspective visual input that is processed via an encoder (shallow ResNet with 3 convolutional blocks; matching the SOTA in DeepMind Lab environment (Espeholt et al., 2018), pretrained and fixed in our experiments). These outputs are linearly mapped to $F=16$ features (FC: fully connected layer), and then sparsified using batch normalization and high thresholding ($\tau=2.43$), such that the percentage of activation ($\sim 2.5\%$) matches the sparse activity of DG granular cells that project to CA3. CA3 is modeled via sequences of neuron activations, each sequence evoked by a distinct DG input feature. The activity of all CA3 neurons are then flattened and linearly mapped to the decoder multilayer perceptron. CA3 is hard coded to isolate the effect of long range integration. The DG and decoder modules are trained.
@@ 2.1 Virtual Environment
We consider an agent navigating a continuous environment with sparse obstacles and uniform visual textures. The environment was simulated using DeepMind Lab (Beattie et al., 2016) (Figure 1D). The wall layouts were kept fixed for repeated episodes unless otherwise stated. The environment was designed such that spatial relations between locations cannot be trivially inferred from the similarity of their corresponding visual features and that it allows multiple possible trajectories from a given location to the target.
In each episode, the agent was initially placed at a random location at least 5 units from the goal. An episode is finished when the agent reaches the goal or a maximum of 7200 frames (with action repeat=8 corresponding to a maximum of 900 action steps, i.e. time steps in the model). For the map displayed in Fig. 1D, the reward was placed near the bottom right corner and then moved to the lower left corner to test generalization.
@@ 2.2 Visual processing
The visual encoder is intended to mimic visual cortical preprocessing, extracting general-purpose visual features for the hippocampus. We pretrained a ResNet encoder (He et al., 2015) in combination with an LSTM core as in Espeholt et al. (2018) and kept it fixed when training our hippocampus model. We also tested a variant using the layer 2 output from ResNet pretrained on Imagenet. The model produces similar performance (Fig. A2).
@@ 2.3 Hippocampus Model
[…]
@@
We model CA3 as a linear RNN shift register that propagates inputs as theta sequences (Fig. 1C and E, CA3) and, for this paper, we keep it fixed to isolate the effects of intrinsic sequence generation from effects potentially induced by recurrent plasticity. The DG provides input $u_{t}\in\mathbb{R}^{F}$ over $F$ features. Each feature is assigned a dedicated prewired sequence of length $\ell$, so the CA3 state is $X_{t}\in\mathbb{R}^{F\ell}$. Motivated by hippocampal theta sequences (Dragoi & Buzsáki, 2006; Foster & Wilson, 2007; Leibold, 2020), we parameterize the total number of sequence units as $\ell=L+(R-1)$, where $L$ is the number of theta cycles and $R$ the number of active units per cycle.
@@
where $I_{F}$ is an identity matrix that has the size of the number of DG features. $S$ and $J$ are the same as the single feature case (For an extended explanation, see Appendix B.2). and the full CA3 dynamics
@@ 2.4 Actor Critic Network
The activity of hippocampal units is fed into a decoder module with two linear/fully connected (FC) layers with ReLU activation. Actions and value are computed as the linear readout from the decoder module. We trained the agent with a standard advantage actor–critic objective (policy-gradient + value-baseline + entropy regularization) as implemented in Sample Factory (Petrenko et al., 2020; Espeholt et al., 2018).
Figure 2: Training performance A Agents with different sequence length $L$ and repeat $R$. Line and shaded area are mean and s.e.m. across 6 random seeds. Success Rate: fraction of episodes where the goal is reached. Len: number of frames to reach the goal. Stable success: the rate of having 100 consecutive successful episodes. Metrics were Gaussian smoothed, $\sigma=6\times 10^{6}$ frames. B The best performing agent with $L=64$ and $R=8$ across all seeds was tested for its transfer learning. New Rew.: new reward location at the lower left corner. new map: a randomly generated map with the same statistics. block path: new walls are added to block paths to the reward, while the previous blocking walls were removed. C and D Performance of agents with different recurrent modules and inputs. C: Sparse input. D: Dense input where batch-normalization and high thresholding was removed. CA3: our CA3 model with $L=64$ and $R=8$. RandRNN: randomly initialized fixed RNN of the same state size. SSM_LegS: fixed SSM HiPPO-LegS from Gu et al. (2020) with the same state size. LSTM: trainable LSTM with matching number of parameters.
@@ 3.1 Behavioral Performance
Training the naive agent ($L=64,R=8$) on one fixed reward location in the maze exhibits robust performance measures after about 350 million frames (Fig. 2A). Thus our model effectively maintains information about sparse inputs in its trajectories of the RNN-like dynamics across time like in a reservoir (Jaeger & Haas, 2004). This view is further supported by our simulations with reduced sequence length showing inferior performance (Fig. 2A). In the most extreme case ($L=1,R=1$) without sequences, essentially a pure feedforward architecture corresponding to a brain with DG output bypassing CA3, the agent did not achieve robust behavior. While agents with reduced sequence lengths ($L=1,16,R=8$) still can express reasonable success rates and trajectory lengths, their behavior is considerably more unstable as evidenced by the low fraction of consecutive successes (Fig. 2A). The agent’s performance is stable within a wide range of $R$ and becomes more sensitive to $R$ when running speed is slower (Fig. A1 and Tab. A3). Transfer learning for new reward location and blocked paths, however, requires only about 50 million frames, and transfer learning to a new map requires about 150 million frames. These indicate that the agent was able to form a generalizing representation of the map and task (Fig. 2B).
@@
Crucially, this advantage is confined to the sparse-input regime. Under dense input, LSTMs perform better—consistent with prior reports (Hessel et al., 2019)—and our CA3 module performs worse than HiPPO-LegS and random RNNs (Fig. 2D). We also tested other variants from Gu et al. (2020) and they performed similar to HiPPO-LegS (Tab. 1, Fig. A4). These results highlight a specific synergy between sparse representations and intrinsic sequence generator.
Figure 3: Evolution of occupancy over the course of learning. Color represents a normalized measure of the amount of time points the agent spent at a location. Mean head directions at a location were visualized in arrows. From epoch 5 onwards, the agent has a preference to reach the bottom right corner (independent of the random starting location) and proceed to the goal (red star) from there. The last panel shows the behavior of fully trained dense LSTM agent.
@@ 3.2 Occupancy
We divided the training into 6 epochs and evaluated the behavior and learned representations at these checkpoints. The agent develops a stable trajectory after 4 to 5 epochs. Between Epoch 4 and Epoch 5, the agent learned to get around the obstacle in the upper part of the right edge and started to spend more time at the lower right corner before reaching the reward site (Fig. 3). The agent appears to develop a strategy of visiting locations with salient input/landmarks and converging from different starting locations to the habituated paths, even after switching to new reward location. This is consistent with the typical strategies employed by animals in familiar enviroments (Gibson et al., 2013). In comparison, the LSTM agent under dense sensory input has fewer converging trajectories before reaching the reward (Fig. 3 and Fig. A11, likely due to the goal information being more readily available in the visual input, thereby implementing a strategy more related to visual search.
@@
Figure 4: A Spatial tuning of DG and CA3 units from epoch 6. Pixel coordinates correspond to environment (black crosses correspond to walls). Each row shows the CA3 units ordered by their positions in the activity sequence. We selected 4 out of the 16 feature sequences for visualization. Spatial Information (SI): bits per time step. B Spatial tuning of randomly selected Decoder layer 1 units from epoch 6. C Spatial tuning of example LSTM output units in the dense LSTM agent. Colormaps are min-max scaled.
Figure 5: A Correlation (color code) of the 16 DG activity maps evoked by the visual features. Throughout training, the DG input to CA3 became increasingly orthogonal. B The spread of CA3 place field measured by entropy. The units are ordered by their positions in a sequence. Each sequence is indicated by a unique color across epochs. C Same as B but for the spatial information of CA3 units.